OpenStudy (anonymous):

solve: ^3√x+8 = 5 my answer is x=117

OpenStudy (anonymous):

is this \[\sqrt[3]{x+8}=5\]?

OpenStudy (anonymous):

get \[x+8=125\] and you are right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cube root x + 8 = 5 Then you have cube root x = -3 Cubing both sides gives x = 27. If the entire x + 8 is the cube root, then you do the same logic, and yes it is x = 117. http:/www.tutorsean.net

OpenStudy (anonymous):

thanks for your video site you're awesome